OpenStudy (anonymous):

TEXASCHIC101, help! =)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

So did you see part c to get started?

OpenStudy (anonymous):

did you type something?

OpenStudy (anonymous):

so are you able to help?

OpenStudy (anonymous):

r u going to help?

OpenStudy (texaschic101):

let me check something slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (29 - 7) / (1 - 0) slope(m) = 22 y = mx + b 7 = 22(0) + b 7 = b you got this right....y intercept is 7 ===================== average rate of change from x = 2 to x = 4 slope = (95 - 51) / (4 - 2) slope = 44/2 slope = 22 (same as average rate of change) ====================== well...the domain is all the x values 227 miles...well the slope is 22 f(x) = 22x + 7 f(10) = 22(10) + 7 f(10) = 227 my guess, and I am not 100% sure, would be : 9 < x < 11 ?? I say this because for him to travel 227 miles, it would take him 10 hrs

OpenStudy (anonymous):

Thank you though I did it myself, i gave you a medal you deserve it! =)

OpenStudy (texaschic101):

sorry that I wasn't much help

OpenStudy (anonymous):

also i wanted help not the answer, but i put something else either way